OpenStudy (raffle_snaffle):

Consider this combination reaction: 2Mg(s)+O2(g)→2MgO(s) ΔH=−1204 kJ What is the enthalpy for the decomposition of 1 mole of MgO(s) into Mg(s) and O2(g)? Consider this combination reaction: What is the enthalpy for the decomposition of 1 mole of into and ? -1204 kJ/mol 602 kJ/mol 1204 kJ/mol -602 kJ/mol

OpenStudy (raffle_snaffle):

it's (a) right? This is a trick question or am I just too dumb to see this? @Photon336

OpenStudy (raffle_snaffle):

@Kainui

OpenStudy (kainui):

I'm confused what the question is, it looks like there are two here and the second one got cut off

OpenStudy (raffle_snaffle):

well A isn't the answer lol I am confused too

OpenStudy (raffle_snaffle):

hold on....

OpenStudy (raffle_snaffle):

Consider this combination reaction: 2Mg(s)+O2(g)→2MgO(s) ΔH=−1204 kJ What is the enthalpy for the decomposition of 1 mole of MgO(s) into Mg(s) and O2(g)?

OpenStudy (raffle_snaffle):

Sorry did a copy and paste, apparently it didn't copy everything

OpenStudy (kainui):

Well it looks like you're asking for what's the reverse reaction of a single mole of MgO. So you'll have to change two things. Right now you lose 1204 kJ of heat to go from left to right. But you now want to go from right to left, so you need to gain 1204 kJ of heat to reverse it. So \(\Delta H = +1204 kJ\) for the reverse... However this isn't all, cause that reaction is for 2 moles of MgO. So now you just divide it by 2. Done!

OpenStudy (raffle_snaffle):

Why do we divide by 2?

OpenStudy (raffle_snaffle):

The ratio is 1:1

OpenStudy (raffle_snaffle):

2:2

OpenStudy (kainui):

Ahhh good question, it's because it takes half as much heat for half as much molecules. The difference is between what you're given here: 2Mg(s)+O2(g)→2MgO(s) ΔH=−1204 kJ and what your reaction is concerned with is a single mole of MgO: Mg(s)+(1/2)O2(g)→MgO(s) ΔH=−602 kJ Everything is reduced by a half

OpenStudy (kainui):

Well wait, you didn't really say anything specific, just Enthalpy not Molar Enthalpy right?

OpenStudy (raffle_snaffle):

Yeah, but I see what you are doing there. You are using least common multiple.

OpenStudy (raffle_snaffle):

mathematically I understand but conceptual having a bit of ahard time seeing it.

OpenStudy (raffle_snaffle):

This heat we are looking at is the amount of heat causing the chemical reaction? It takes 602kJ to break the bonds?

OpenStudy (kainui):

Really what I'm saying is 2MgO takes 1204 kJ of heat to decompose so if you have half as much, so that means 1MgO then it must take half as much heat to decompose, 602 kJ. Imagine you have two pots of water and one has twice as much water than the other, then it takes twice as much heat to get it to boil. Same kinda idea.

OpenStudy (raffle_snaffle):

Oh okay lol thanks that makes more sense!

OpenStudy (kainui):

Cool :P